OpenStudy (anonymous):

what is this quotient in simplest form? x+2/x-1divided by x+4/x^2+4x-5

OpenStudy (whpalmer4):

\[\frac{\frac{(x+2)}{(x-1)}}{\frac{(x+4)}{(x^2+4x-5)}} = \frac{(x+2)}{(x-1)}*\frac{(x^2+4x-5)}{(x+4)}\] because dividing by a fraction is equivalent to multiplying by its reciprocal Now factor (x^2+4x-5) and cancel out any matching expressions.